OpenStudy (anonymous):

sqrt36x^3/y^2 = 6x/y True or False?

OpenStudy (lgbasallote):

\[\large \sqrt{\frac{36x^3}{y^2}} \implies \sqrt{\frac{6^2 \times x^2 \times x}{y^2}} \implies\frac{6x}{y}\sqrt x\] does that help?

OpenStudy (anonymous):

yeah thanks :)

OpenStudy (lgbasallote):

welcome